Find the general integral of the equation
(x − y) p + (y − x − z) q = z
and a particular solution through the circle z =1, x^2 + y^2 =1
1
Expert's answer
2020-04-01T16:20:39-0400

"(x-y)p+(y-x-z)q=z\\
z=1, x^2+y^2=1"

Write the system in the symmetric form

"\frac{dx}{x-y}=\frac{dy}{y-x-z}=\frac{dz}{z}\\
i)\frac{dx}{x-y}=\frac{dy+dz}{y-x} \\
-dx=dy+dz\\
dx+dy+dz=0\\
d(x+y+z)=0\\
x+y+z=c_1\\
ii) x+z=c_1-y\\
\frac{dz}{z}=\frac{dy}{y-x-z}\\
\frac{dz}{z}=\frac{dy}{y-(x+z)}\\
\frac{dz}{z}=\frac{dy}{y-(c_1-y)}\\
\frac{dz}{z}=\frac{dy}{2y-c_1}\\
\int \frac{dz}{z}=\int\frac{dy}{2y-c_1}\\
\ln|z|+\frac{1}{2}\ln|c_2|=\frac{1}{2}\ln|2y-c_1|\\
c_2z^2=2y-c_1\\
c_2z^2=2y-x-y-z\\
\frac{y-x-z}{z^2}=c_2"

General solution is

"F(x+y+z,\frac{y-x-z}{z^2})=0"

Find particula solution

"\left\{\begin{matrix}
 z=1 \\
 x^2+y^2=1\\
x+y+z=c_1\\
\frac{y-x-z}{z^2}=c_2
\end{matrix}\right.\\
\left\{\begin{matrix}
 x^2+y^2=1\\
x+y+1=c_1\\
y-x-1=c_2
\end{matrix}\right.\\
\left\{\begin{matrix}
 x^2+y^2=1\\
x^2+y^2+2xy=(c_1-1)^2\\
x^2+y^2-2xy=(c_2+1)^2
\end{matrix}\right.\\
\left\{\begin{matrix}
 1+2xy=(c_1-1)^2\\
1-2xy=(c_2+1)^2
\end{matrix}\right.\\
\left\{\begin{matrix}
 2xy=(c_1-1)^2-1\\
2xy=1-(c_2+1)^2
\end{matrix}\right.\\
(c_1-1)^2-1=1-(c_2+1)^2\\
(c_1-1)^2+(c_2+1)^2=2"

So particula solution is

"(x+y+z-1)^2+(\frac{y-x-z}{z^2}+1)^2=2"